OpenStudy (anonymous):

factor m from the expression: 5mn=3m

OpenStudy (anonymous):

m=0 is that correct?

OpenStudy (anonymous):

well i got m = 0 and n =-8 im not sure what i should put down as my aswer lol

OpenStudy (anonymous):

so if u got m=0 n i did too then i guess thats the answer

OpenStudy (anonymous):

I think u will have to use the m=0, because u were asked to factor out m from the expression

OpenStudy (anonymous):

ive been awake for to long lol

OpenStudy (anonymous):

Oh then you will be very tired then. but what were you doing all along?

OpenStudy (unklerhaukus):

\[5mn=3m\]\[\quad5n=3\]

OpenStudy (anonymous):

How? why did you cancel out the m's ?

OpenStudy (unklerhaukus):

they were common

OpenStudy (anonymous):

but don't you think in this case you are suppose to make m the subject?

OpenStudy (unklerhaukus):

the question asks us to factor m from the expression I factored m right out of the expression

OpenStudy (anonymous):

ok, think av got ur idea!

OpenStudy (anonymous):

unkle u rock bro

OpenStudy (anonymous):

worked night shift, got class in a couple hours and im cathcing up on homework

OpenStudy (anonymous):

Uncle, take a luk at the question again, u wer asked to 'factor m from' the expresion, so m must not be missin from it but rather ,it must stand out in the expression

OpenStudy (anonymous):

so is it m = 0 or n=-8m?

OpenStudy (anonymous):

In that case , 5mn=3m => 5mn-3m=0 m(5n-3)=0 m=0/(5n-3) m=0

OpenStudy (anonymous):

you guys seem to be having fun with this one

OpenStudy (anonymous):

sure! u'v got help us out.

OpenStudy (anonymous):

haha well factoring the m out is still factoring it and whateve process we use we end up with m=0

OpenStudy (anonymous):

well first i came to n=-8m and that is simplified to 0 anyway

OpenStudy (anonymous):

so i like the process unkle used cause its simple that way